# Syntax-Guided Procedural Synthesis of Molecules

### v1

@@ 3.3.1 Outer-Loop: Syntax Tree Structure Optimization
We simulate a Markov Process on $\mathcal{T}$, the set of all binary trees for discovering tree structures that maximize similarity between $M$ and the program’s output. The details for how we bootstrap and apply $\mathcal{T}$ is in App A. We adopt the Metropolis-Hastings algorithm with proposal distribution $J(T_{1},T_{2})\propto\exp(-\lambda\cdot\text{dist}(T_{1},T_{2}))$. The scoring function is $\pi(T)\propto\exp(-\beta\cdot\text{dist}(M,\text{exec}(F(M,T))))$ where $\lambda,\beta$ are parameters to tradeoff exploration with exploitation. In other words, we use the inner-loop to score candidates in $\mathcal{T}_{k}$. The optional outer loop benefits from the amortized nature of the inner loop. This highlights how our multi-step factorization decouples the syntactic and semantic components of a complete solution. Next, we discuss our amortized learning strategy for each step.
3.3.2 Inner-loop: Inference of Tree Semantics
We formulate the conditional derivation of $G_{\mathcal{P}}$ after the syntax tree is fixed as a finite horizon Markov decision process (MDP). To bridge $\mathcal{T}$ and $\mathcal{P}$, we introduce an intermediate space of partial programs $\mathcal{T^{\prime}}$. $\mathcal{T^{\prime}}$ consists of all possible partial programs that arise from the following modifications to a $T\in\mathcal{T}$:
  1. 1.
    Prepend a new root node as the parent of the root node of $T$.
  2. 2.
    For the root node, attribute it with any $x\in\mathcal{X}$.
  3. 3.
    Attribute a subset of $T$ with reactions from $\mathcal{R}$.
  4. 4.
    For each leaf node of $T$, attach a left child, and optionally attach a right child.
  5. 5.
    For the newly attached leaf nodes, attribute a subset of them with building blocks from $\mathcal{B}$.
Intuitively, $\mathcal{T}^{\prime}$ is the space of all partially filled in trees in $\mathcal{T}$, with the caveats of adding a root node attributed with a fingerprint describing a hypothetical molecule and adding leaf nodes representing building blocks.
State Space: The state space $S\subset\mathcal{T^{\prime}}$ is the set of all partially filled in trees satisfying topological order, i.e., $S=\{T\in\mathcal{T^{\prime}}\mid\text{$n$ is filled}\Rightarrow\text{parent($n$) is filled, for all }n\in T\}$.
Action Space: The actions $A_{s}$ for a given state available are to fill in a frontier node, i.e., $\{n\in T\mid\text{parent($n$) is filled}\}$ with a reaction from $\mathcal{R}$ if $n$ is a reaction and a building block from $\mathcal{B}$ otherwise.
@@ 3.3.2 Inner-loop: Inference of Tree Semantics
Policy Network: We parameterize policy networks $G_{R;\Phi}\colon\mathcal{T^{\prime}}\rightarrow\mathcal{R^{*}}$ and $G_{B;\Omega}\colon\mathcal{T^{\prime}}\rightarrow\mathcal{B^{*}}$ by separate graph neural networks, both taking as input a partial program $T^{\prime}$. The network $G_{R;\Phi}$ predicts reactions at the unfilled reaction nodes of $T^{\prime}$, while $G_{B;\Omega}$ predicts embeddings at the unfilled leaf nodes. Since $|\mathcal{B}|\gg|\mathcal{R}|$, $G_{B;\Omega}$ instead predicts 256-dimensional continuous embeddings, from which we retrieve the building block whose 256-dimensional Morgan fingerprint is closest.
@@
Training: We train $\Phi,\Omega$ using supervised policy learning. The key to this approach is the dataset used for training, $D_{\text{pretrain}}$. We construct $D_{\text{pretrain}}$ as follows:
For each (program $P^{(i)}$, building blocks $B^{(i)}$, molecule $M^{(i)}$) in our synthetic dataset,
  1. 1.
    Construct $T^{\prime(i)}$ by prepending a new root node attributed with the fingerprint of $M^{(i)}$ and attaching leaf nodes $L$ corresponding to $B^{(i)}$.
  2. 2.
    For each $\texttt{mask}\in\{0,1\}^{|T^{\prime(i)}|}$ such that $\texttt{mask}[\text{root}(T^{\prime(i)})]$ and $\texttt{mask}[i]\Rightarrow\texttt{mask}[\text{parent}(i)]$ for all $i\in T^{\prime(i)}\setminus\{\text{root}(T^{\prime(i)})\}$,
    1. 1.
      Obtain the frontier nodes $\text{Fr}_{\texttt{mask}}\leftarrow\{n\in T^{\prime(i)}\mid\texttt{!mask}[n]\cap\texttt{mask}[\mathrm{parent}(n)]\}$
    2. 2.
      Initialize $(X^{(i)},y^{(i)})$ with all the node features and labels
    3. 3.
[…]
    4. 4.
      Update $D_{\text{pretrain}}\leftarrow D_{\text{pretrain}}\cup\{(T^{\prime(i)},X^{(i)},y^{(i)})\}$
Additional details are in App E.
Decoding: Our decoding algorithm is designed to align with the pretraining strategy, shown in Figure 3. Instead of performing search over an indefinite horizon, we restrict the set of horizon structures to $\mathcal{T}_{k}$, i.e., those present in $D_{\text{pretrain}}$. At test time, we either: (a) use $Q_{\Theta}$ to recognize $T\in\mathcal{T}_{k}$ and initialize the decoding process, or (b) use the algorithm described in Section 3.3.1 to simulate a Markov Chain over $\mathcal{T}_{k}$ and refine the skeleton over time.
@@
Figure 3: Illustration of our decoding scheme $F_{\phi}(M)$: (Left) The input is a syntax skeleton $T$ and Morgan fingerprint $X$; (Middle) Decode once for every topological ordering of the tree, tracking all partial programs with a stack; (Right) Execute all decoded programs, then returning the closest analog which minimizes distance to $M$.
3.4 Bilevel Syntax-Semantic Synthesizable Molecular Design
The task of synthesizable molecule design is to solve $\operatorname*{arg\,min}_{(P,B)\in\mathcal{P}\times 2^{\mathcal{B}}}f(P(B))$ for a property oracle $f$ of interest. Given the learned parameters from synthetic planning $\Phi,\Omega$, we apply the inner-loop procedure $F_{\Phi,\Omega}(M,T)\colon\mathcal{X}\times\mathcal{T}_{k}\rightarrow\mathcal{P}\times 2^{\mathcal{B}}$ as a surrogate, casting the problem as $\operatorname*{arg\,min}_{(X,T)\in\mathcal{X}\times\mathcal{T}_{k}}f(\text{exec}(F_{\Phi,\Omega}(X,T)))$. This two-step factorization of the design space enables joint optimization over both the semantics and syntax of a solution.
We approach this problem with a Genetic Algorithm (GA) over the joint design space $\mathcal{X}$ and $\mathcal{T}_{k}$. The seed population is obtained by sampling random bit vectors and using $Q_{\Theta}$ to obtain their corresponding skeletons. To generate children $(X,T)$ from two parents $(X_{1},T_{1})$ and $(X_{2},T_{2})$, we combine semantic crossover with syntactical mutation, consistent with our bilevel approach for analog generation:
@@ 3.4 Bilevel Syntax-Semantic Synthesizable Molecular Design
Semantic Evolution: We generate $X$ by combining bits from both $X_{1}$ and $X_{2}$ and possibly mutating a small number of bits of the crossover result. We then set $T=Q_{\Theta}(X)$.
Syntactic Mutation: We set $X=X_{1}$ and apply edit(s) to $T_{1}$ to obtain $T^{\prime}$. We set $T\leftarrow\text{select}(T^{\prime},T_{1})$ with the selection criteria determined by an exploration-exploitation annealing schedule.
@@
Intuitively, Semantic Evolution optimizes for chemical semantics by combining existing ones from the mating pool, while Syntactic Mutation explores syntactic analogs of specific individuals of interest. Notably, our implementation borrows ideas from simulated annealing [34] to facilitate global diversity during the early iterations of GA and local exploitation during the later iterations of GA. Offspring are generated through a combination of both strategies. Each child is then decoded into a SMILES string using the surrogate and given a fitness score under the property oracle. We then reassign the children’s fingerprints after decoding, i.e., $X\leftarrow\text{FP}(\text{exec}(F_{\Phi,\Omega}(X,T))))$ to reflect our analog exploration strategy. The top scoring unique individuals between are retained into the next generation (i.e., elitist selection). Further details and hyperparameters are given in App F.
@@ 4.1.1 Data Generation
We use 91 reaction templates from [27, 6] representative of common synthetic reactions. They consist primarily of ring formation and fusing reactions but also peripheral modifications. We use 147,505 purchaseable building blocks from Enamine Building Blocks. We follow the same steps as [21] to generate 600,000 synthetic trees. After filtering by QED, we obtain 227,808 synthetic trees (136,684 for training, 45,563 for validation, and 45,561 for test). We then pre-process these trees into programs, constructing our dataset $D_{\text{train}},D_{\text{valid}}$ and $D_{\text{test}}$. We bootstrap our set of syntactic templates, $\hat{\mathcal{T}}$ based on those observed in $D_{\text{train}}$, resulting in 1117 syntactic skeleton classes. Additional statistics on $\hat{\mathcal{T}}$ and insights on its coverage are given in App A.
@@ 4.1.2 Baselines
We evaluate against all 25 molecular optimization methods benchmarked in the large-scale experimental study [22]. These methods are divided into three categories, based on the molecule representation: string, graph, or synthetic trees. Synthesis methods restricts the design space to only products of robust template-based reactions, so for fair comparison, so we also report intra-category rankings. We also report the SA Score [18] of the optimized molecules for all methods, both to cross-verify the synthesizability of synthesis-based methods and to investigate the performance trade-off imposed by constraining for template-compatible synthesizability.
@@ 4.1.3 Evaluation Metrics
Synthesizable Analog Generation: We evaluate the ability to generate a diverse set of structural analogs to a given input molecule using Average Similarity (as measured by Tanimoto distance to the input) and Internal Diversity (average pairwise Tanimoto distance). We also include the Recovery Rate (RR), whether the most similar analog reconstructs the target, and SA Score [18] as a commonly-used heuristic for synthetic accessibility.
Synthesizable Molecule Design: We evaluate the ability to optimize against calls to 15 Oracle functions [29] relevant to real-world drug discovery. In addition to Top-k average score, we particularly focus on sample efficiency (Top k AUC), as described in [22]. We now describe the different Oracle categories:
@@
    Structural profiles (Median1/Median2/Rediscovery): These primarily focus on maximizing structural similarity to multiple molecules, useful for designing molecules fitting a more multifaceted structural profile [4]. The rediscovery oracle focuses on hit expansion around a specific drug.
@@
    Multi-property objectives (Osimertinib & 6 others): These use real drugs as a basis for optimizing additional pharmacological properties, mimicking how real-world drug discovery works.
@@
    Docking Simulations ($M^{\text{pro}}$, DRD3): We also add two Docking simulation Oracles against $M^{\text{pro}}$, the main protease of SARS-Cov-2 and DRD3, which has its own leaderboard, with a particular focus on sample efficiency.
Table 1: We generate 5 unique analog molecules conditioned on an input molecule $M$ and sort them by decreasing similarity to $M$. For SynNet, we follow their beam search strategy and produce analogs using the top 5 beams. For Ours (Skeleton), we sample the top 5 syntactic templates from $Q_{\Theta}$. Then, we evaluate how similar, diverse, and structurally simple the first $k$ molecules are. The best method is bolded. For ChEMBL, the second best method is underlined.
Dataset Method RR $\uparrow$ Avg. Sim. $\uparrow$ SA Score $\downarrow$ Diversity $\uparrow$
$k=1$ $k=3$ $k=5$ $k=1$ $k=3$ $k=5$ $k=3$ $k=5$
Test Set SynNet 46.3% 0.766, 0.622, 0.566 3.108, 3.057, 3.035 0.525, 0.584
Ours (Skeleton) 52.3% 0.799, 0.588, 0.548 3.075, 2.895, 2.856 0.609, 0.653
@@
4.2 Results
4.2.1 Synthesizable Analog Generation
@@
In Table 1, we see our method outperforming SynNet across both dimensions of similarity (how “analog” compounds are) and diversity (how different the compounds are). Additionally, our method achieves lower SA Score, which is a proxy for synthetic accessibility that rewards simpler molecules. Guided by a set of simple yet expressive syntactic templates, our model simultaneously produces more diverse and structurally simple molecules without sacrificing one for the other. Additionally, our policy network is well-suited to navigate these simple yet horizon structures, enabling a $6\%$ higher reconstruction accuracy after training on the same dataset. Combining these three dimensions, we can conclude our method is the superior one for the task of synthesizable analog generation. To better understand which design choices are responsible for the performance, we provided a comprehensive analysis of the policy network in App E. We begin in App E.3 by elaborating on the main distinction of our method vs existing works, highlight the novelty of our formulation, and motivate an auxiliary training task that takes inspiration from cutting-edge ideas in inductive program synthesis. We then perform several key ablations in App E.4, using concrete examples to highlight success and failure cases. Lastly, we perform a step-by-step walkthrough of our decoding algorithm in App E.6, visualizing the evolution of attention weights to showcase the full-horizon awareness of our surrogate and the dynamic incorporation of new information. These analyses shed insights into why our surrogate works, and points to future extensions to make it even better.
4.2.2 Synthesizable Molecule Design
Table 2: We limit to 10000 Oracle calls, then compute Top $k$ and Top $k$ AUC following the settings of [22]. We also include synthetic accessibility scores [18]. The best method per column is bolded, and the best synthesis-based method is underlined. Each Oracle’s output is normalized to $[0,1]$. We compute the Average across all 13 Oracles here. See App. G for the full results and experiment details.
Average (13 properties)
Top 1 Top 10 Top 100
category Score SA AUC Score SA AUC Score SA AUC
synnet synthesis 0.603 (9|3) 3.074 (7|4) 0.577 (7|2) 0.578 (9|3) 3.075 (6|4) 0.545 (7|2) 0.527 (10|3) 3.094 (7|4) 0.48 (7|2)
pasithea string 0.418 (24|10) 3.783 (15|7) 0.404 (23|9) 0.338 (24|10) 3.66 (12|5) 0.326 (24|10) 0.249 (24|10) 3.835 (13|6) 0.24 (24|10)
dog_ae synthesis 0.508 (17|4) 2.845 (4|3) 0.501 (14|4) 0.46 (18|4) 2.857 (3|3) 0.45 (14|4) 0.353 (19|4) 2.809 (3|3) 0.339 (16|4)
smiles_vae_bo string 0.486 (21|9) 3.18 (8|2) 0.456 (20|8) 0.422 (21|9) 3.084 (7|2) 0.376 (21|8) 0.323 (20|8) 3.055 (4|1) 0.284 (20|8)
jt_vae_bo graph 0.469 (22|7) 3.571 (11|1) 0.453 (21|7) 0.388 (23|8) 3.5 (10|1) 0.371 (22|8) 0.293 (23|8) 3.559 (10|1) 0.278 (23|8)
moldqn graph 0.238 (26|10) 5.4 (25|10) 0.209 (26|10) 0.213 (26|10) 5.604 (25|10) 0.187 (26|10) 0.177 (26|10) 5.69 (25|10) 0.151 (26|10)
mars graph 0.539 (15|5) 4.148 (21|8) 0.508 (13|4) 0.507 (14|5) 4.232 (21|8) 0.47 (12|4) 0.463 (14|5) 4.436 (22|9) 0.41 (13|5)
selfies_lstm_hc string 0.579 (11|5) 3.617 (12|5) 0.49 (16|6) 0.539 (12|6) 3.743 (14|6) 0.431 (16|6) 0.485 (13|6) 3.82 (12|5) 0.351 (15|6)
gp_bo graph 0.662 (7|2) 3.981 (18|5) 0.599 (5|2) 0.642 (6|2) 3.954 (16|3) 0.57 (5|2) 0.617 (6|2) 4.054 (17|4) 0.524 (6|2)
smiles_ga string 0.555 (12|6) 5.107 (23|8) 0.519 (11|5) 0.548 (11|5) 5.422 (23|8) 0.503 (10|5) 0.537 (9|5) 5.578 (23|8) 0.479 (8|4)
mimosa graph 0.551 (13|4) 4.17 (22|9) 0.499 (15|5) 0.538 (13|4) 4.3 (22|9) 0.463 (13|5) 0.515 (12|4) 4.378 (21|8) 0.417 (12|4)
reinvent string 0.711 (2|1) 3.352 (9|3) 0.633 (2|1) 0.697 (2|1) 3.415 (9|3) 0.607 (2|1) 0.685 (1|1) 3.48 (9|3) 0.573 (1|1)
smiles_lstm_hc string 0.695 (3|2) 3.016 (5|1) 0.599 (5|3) 0.667 (5|3) 3.036 (5|1) 0.544 (8|4) 0.622 (5|3) 3.055 (4|1) 0.462 (9|5)
selfies_vae_bo string 0.502 (18|7) 3.423 (10|4) 0.465 (17|7) 0.428 (19|8) 3.522 (11|4) 0.383 (19|7) 0.318 (22|9) 3.628 (11|4) 0.281 (22|9)
dog_gen synthesis 0.663 (6|2) 2.766 (3|2) 0.562 (9|3) 0.634 (7|2) 2.793 (2|2) 0.511 (9|3) 0.591 (8|2) 2.803 (2|2) 0.424 (11|3)
stoned string 0.613 (8|4) 5.364 (24|9) 0.568 (8|4) 0.609 (8|4) 5.55 (24|9) 0.555 (6|3) 0.599 (7|4) 5.667 (24|9) 0.529 (4|2)
gflownet graph 0.495 (19|6) 4.069 (19|6) 0.461 (19|6) 0.461 (17|6) 4.05 (19|6) 0.419 (17|6) 0.403 (16|6) 4.17 (19|6) 0.359 (14|6)
reinvent_selfies string 0.693 (5|3) 3.73 (13|6) 0.608 (4|2) 0.682 (3|2) 3.791 (15|7) 0.578 (4|2) 0.654 (3|2) 3.856 (15|7) 0.528 (5|3)
graph_mcts graph 0.37 (25|9) 3.745 (14|2) 0.332 (25|9) 0.317 (25|9) 3.732 (13|2) 0.28 (25|9) 0.246 (25|9) 3.849 (14|2) 0.213 (25|9)
dst graph 0.584 (10|3) 4.125 (20|7) 0.522 (10|3) 0.555 (10|3) 4.146 (20|7) 0.479 (11|3) 0.52 (11|3) 4.29 (20|7) 0.426 (10|3)
selfies_ga string 0.495 (19|8) 5.693 (26|10) 0.358 (24|10) 0.48 (15|7) 5.709 (26|10) 0.337 (23|9) 0.455 (15|7) 5.861 (26|10) 0.306 (19|7)
gflownet_al graph 0.463 (23|8) 3.829 (16|3) 0.434 (22|8) 0.417 (22|7) 4.005 (18|5) 0.387 (18|7) 0.354 (18|7) 4.153 (18|5) 0.32 (18|7)
screening N/A 0.52 (16|2) 3.042 (6|2) 0.464 (18|2) 0.426 (20|2) 3.097 (8|2) 0.377 (20|2) 0.322 (21|2) 3.064 (6|1) 0.284 (20|2)
mol_pal N/A 0.548 (14|1) 2.64 (2|1) 0.517 (12|1) 0.472 (16|1) 3.018 (4|1) 0.444 (15|1) 0.366 (17|1) 3.105 (8|2) 0.339 (16|1)
graph_ga graph 0.716 (1|1) 3.835 (17|4) 0.631 (3|1) 0.701 (1|1) 3.982 (17|4) 0.601 (3|1) 0.676 (2|1) 4.018 (16|3) 0.553 (3|1)

### v2

@@ 3.3.1 Outer loop: Syntax Tree Structure Optimization
We simulate a Markov process on ${\mathcal{T}}$ for discovering skeletons whose decoding will maximize the similarity between $M$ and the program’s output (Figure 2 (Left)). The details for how we bootstrap and apply ${\mathcal{T}}$ is in App. A. We adopt the Metropolis-Hastings algorithm with proposal distribution $q(T\,|\,T_{0})\propto\exp(-\lambda\,d_{{\mathcal{T}}}(T,T_{0}))$ and scoring function $f(T)\propto\exp(-\beta\,d_{\mathcal{M}}(M,F(M,T)))$, where $\lambda$ and $\beta$ are parameters that trade-off exploration with exploitation, $d_{{\mathcal{T}}}$ is the tree edit distance, and $d_{\mathcal{M}}$ is the Tanimoto distance. In other words, we use the inner loop to score candidates in ${\mathcal{T}}$.
3.3.2 Inner loop: Inference of Tree Semantics
We now formulate syntax-conditioned derivations under $G_{{\mathcal{P}}}$ as a finite-horizon MDP.
State space: To bridge ${\mathcal{T}}$ and ${\mathcal{P}}\times{\mathcal{B}}^{*}$, we introduce an intermediate state space of partial programs $\partial{\mathcal{P}}$ consisting of all possible partial programs arising from the following modifications to any syntax tree $T\in{\mathcal{T}}$: (1) prepend a new root node of $T$, and attribute it with a fingerprint from ${\mathcal{X}}$, or (2) attribute some internal (resp. leaf) nodes of $T$ with elements of $\mathcal{R}$ (resp. ${\mathcal{B}}$). We further require that if a node in $T$ is filled, then so is its parent. Intuitively, $\partial{\mathcal{P}}$ comprises all partially filled-in trees in ${\mathcal{T}}$ obeying topological order (with an added root node attached to a molecular fingerprint).
Action space: At a state $S\in\partial{\mathcal{P}}$, the actions are to attribute any frontier node, i.e., unfilled nodes whose parents are filled, with an item from $\mathcal{R}$ (resp. ${\mathcal{B}}$) if the node is internal (resp. a leaf).
@@ 3.3.2 Inner loop: Inference of Tree Semantics
Policy network: We parameterize the policies $\pi_{\mathcal{R}}\colon\partial{\mathcal{P}}\rightarrow\mathcal{R^{*}}$ and $\pi_{{\mathcal{B}}}\colon\partial{\mathcal{P}}\rightarrow\mathcal{B^{*}}$ with separate graph neural networks. Given a partial program $S$ as input, the former predicts the reactions that should be attributed to internal frontier nodes as an $|\mathcal{R}|$-label classification problem, while the latter does so for building blocks and leaf nodes. However, $|{\mathcal{B}}|\gg|\mathcal{R}|$ is very large, so $\pi_{{\mathcal{B}}}$ instead predicts a 256-dimensional embedding at each nodes which implicitly specifies the building block whose 256-bit Morgan fingerprint is closest.
@@
Training: We train both policies using supervised policy learning. Key to this approach is the dataset used for training constructed using Algorithm 1. See App. F for further details.
Algorithm 1 Construction of training dataset.
1:A synthetic dataset ${\mathcal{D}}_{0}\subseteq{\mathcal{P}}\times{\mathcal{B}}^{*}$ of programs (Section 4.1.1).
2:${\mathcal{D}}\leftarrow\varnothing$
3:for each $(P,B)\in{\mathcal{D}}_{0}$ do
4:     Turn $(P,B)$ into a fully-filled program $T\in\partial{\mathcal{P}}$ whose root is attributed with $\text{FP}(P,B)$.
5:     for each $\Lambda\in 2^{T}$ containing the root and closed under $\text{parent}(\cdot)$  do
6:         $\text{Frontier}(\Lambda)\leftarrow\{i\in T\mid i\notin\Lambda\text{ and }\text{parent}(i)\in\Lambda\}$
7:         Populate node features ${\bm{H}}$ and labels ${\bm{Y}}$ based on $P$ and $B$ (App. F)
8:         for $i\in T-\Lambda$ do
9:              Mask the feature in ${\bm{H}}$ corresponding to node $i$
10:         for $i\in T-\text{Frontier}(\Lambda)$ do
11:              Mask the label in ${\bm{Y}}$ corresponding to node $i$
12:         ${\mathcal{D}}\leftarrow{\mathcal{D}}\cup\{(T,{\bm{H}},{\bm{Y}})\}$      return ${\mathcal{D}}$
@@
Figure 3: Illustration of our decoding scheme $F$: (Left) The input is a Morgan fingerprint ${\bm{x}}$ and syntax skeleton $T$; (Middle) Decode once for every topological ordering of the tree, tracking all partial programs with a stack; (Right) Execute all decoded programs, then returning the closest analog which minimizes distance to ${\bm{x}}$.
3.4 Bi-level Syntax-Semantic Synthesizable Molecular Design
The task of synthesizable molecule design is to find a program $P$ and building blocks $B$ whose $M_{*}$ maximizes a property of interest. Given the learned policies from synthetic planning, we apply the inner loop procedure $F\colon{\mathcal{X}}\times{\mathcal{T}}\rightarrow{\mathcal{P}}\times{\mathcal{B}}^{*}$ as a surrogate, casting the optimization problem as one over the joint design space ${\mathcal{X}}\times{\mathcal{T}}$. We approach this problem with a genetic algorithm (GA) over fingerprint space ${\mathcal{X}}$ that leverages this extra dimension of syntax ${\mathcal{T}}$ (Figure 2 (Right)). The seed population is obtained by sampling random bit vectors. To generate a child ${\bm{x}}$ from two parents ${\bm{x}}_{1}$ and ${\bm{x}}_{2}$, we combine semantic crossover with syntactical mutation, reminiscent of our bi-level approach for analog generation:
@@ 3.4 Bi-level Syntax-Semantic Synthesizable Molecular Design
Semantic crossover: We first generate ${\bm{x}}_{*}$ by combining bits from both ${\bm{x}}_{1}$ and ${\bm{x}}_{2}$ and possibly mutating a small number of bits of the result.
Syntactic mutation: We set $T=\tau({\bm{x}}_{*})$ and apply random edit(s) to obtain a perturbed tree $T^{\prime}$. Together, $({\bm{x}}_{*},T)$ and $({\bm{x}}_{*},T^{\prime})$ form a sibling pool. Applying the surrogate $F$ to each child gives two sibling SMILES that are turned into two sibling fingerprints. We fit a Gaussian process on past individuals and select the sibling with the highest expected improvement as the favoured child ${\bm{x}}$.
@@
Intuitively, semantic crossover optimizes for chemical semantics by combining existing ones from the mating pool, while syntactic mutation explores syntactic analogs of individuals of interest. Each child is then given a fitness score under the property oracle, and the top scoring unique individuals are retained into the next generation (i.e., elitist selection). Further details and hyperparameters are given in App. G.
@@ 4.1.1 Data Generation
We use 91 reaction templates from Hartenfeller et al. (2011); Button et al. (2019) representative of common synthetic reactions. They consist primarily of ring formation and fusing reactions but also peripheral modifications. Starting from 147,505 purchaseable compounds from Enamine Building Blocks, we follow the same steps as Gao et al. (2021) to generate 600,000 synthetic trees. Filtering by $\text{QED}>0.5$ of the product molecules leaves 227,808 synthetic trees (136,684 for training, 45,563 for validation, and 45,561 for testing), which are then preprocessed into programs to construct our final datasets. We bootstrap our set of syntactic templates based on those observed in the training set, resulting in 1117 syntactic skeleton classes. Additional statistics on these syntactic templates and insights on their coverage are given in App. A. Further analyses of the structure-property relationship and a detailed case study are given in App. C.
@@ 4.1.2 Baselines
We evaluate against all 25 molecular optimization methods benchmarked in the large-scale study by Gao et al. (2022). Most of these are divided into three categories based on the molecular representation used: string, graph, or synthesis trees. Synthesis methods restrict the design space to only products of robust template-based reactions, so for fair comparison, we also consider intra-category rankings. We report the synthetic accessibility (SA) score (Ertl & Schuffenhauer, 2009) of the optimized molecules to cross-verify the synthesizability of synthesis-based methods as well a investigate the performance trade-off imposed by constraining for template-compatible synthesizability.
4.1.3 Computational Efficiency
Constructing $\mathcal{D}$. Alg. 1 suggests $|\mathcal{D}|=O(2^{\max_{\mathcal{T}}|T|}|\mathcal{D}_{0}|)$ because we compute all the (closed under $\text{parent}(\cdot)$) masks per $P\in\mathcal{D}_{0}$. However, we don’t need to explicitly store $\mathcal{D}$. We only need to precompute the (closed under $\text{parent}(\cdot)$) masks for each $T\in\hat{\mathcal{T}_{k}}$ ($k$ fixed and small), so the running time is $O(\sum_{T\in\hat{\mathcal{T}_{k}}}2^{|T|})$, independent of $\mathcal{D}_{0}$. Besides, the actual number of masks is far smaller than $2^{|T|}$, and high $|T|$ is less represented in $\mathcal{D}_{0}$, since large programs are less likely to be sampled, so in practice $|\mathcal{D}|$ is much smaller than $2^{\max_{\mathcal{T}}|T|}|\mathcal{D}_{0}|$. Detailed statistics are given in App A.
Training with $\mathcal{D}$. During training, we flat-index into $\mathcal{D}_{0}$ and their precomputed masks to perform a pass over $\mathcal{D}$. Despite the larger dataset, we find the total training steps to actually be comparable with Gao et al. (2021). To scale to larger $k$, we propose a stratified sampling strategy that does only a $O(\mathcal{D}_{0})$ pass per epoch, has positive support over $\mathcal{D}$, and represents each $(P,B)\in\mathcal{D}_{0}$ equally. The idea is to sample a constant number of masks $C$ per $P,B\in\mathcal{D}_{0}$, and re-sample each epoch. We show this has complexity linear to the size of $\mathcal{D}_{0}$ per epoch but empirically converges in much fewer training steps than SynNet for $C=1$. Details are in App F.7. We further include an ablation study on the downstream task performance of this simplified training strategy in App F.5.
@@ 4.1.4 Evaluation Metrics
Synthesizable analog generation. We evaluate the ability to generate a diverse set of structural analogs to a given input molecule using the Recovery Rate (RR, whether the most similar analog is exactly the target), Average Similarity (as measured by Tanimoto distance to the input), SA score, and Internal Diversity (average pairwise Tanimoto distance).
Synthesizable molecule design. We evaluate our method’s ability to optimize 15 oracle functions (Huang et al., 2022) relevant to real-world drug discovery:
@@
    Structural profiles (Median1, Median2, Rediscovery) that primarily focus on maximizing structural similarity to multiple molecules, which is useful for designing molecules fitting a more multifaceted structural profile (Brown et al., 2004). The rediscovery oracle focuses on hit expansion around a specific drug.
@@
    Multi-property objectives (Osimertinib, 6 others) that use real drugs as a basis for optimizing additional pharmacological properties, mimicking real-world drug discovery.
@@
    Docking Simulations (Mpro, DRD3) against Mpro, the main protease of SARS-Cov-2, and DRD3, which has its own leaderboard with a particular focus on sample efficiency.
In addition to the average score of the top $k$ molecules, we particularly focus on sample efficiency, i.e., the top-$k$ AUC as described in Gao et al. (2022).
Table 1: We generate 5 unique analog molecules conditioned on an input molecule $M$ and sort them by decreasing similarity to $M$. For SynNet, we follow their beam search strategy and produce analogs using the top 5 beams. For Ours ($\tau$), we sample the top 5 syntactic templates from $\tau{}$. Ours ($\tau$, rev) is the same except we use a bottom-up decoding process, and it is included as an ablation for Section 4.4. Then, we evaluate how similar, diverse, and structurally simple the first $k$ molecules are. The best method is bolded. For three-way comparisons, the second best method is underlined.
Avg. Sim. $\uparrow$ SA $\downarrow$ Diversity $\uparrow$
Dataset Method RR $\uparrow$ Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 Top-3 Top-5
Train Set Ours ($\tau$, rev) 79.3 % 0.923 0.632 0.569 3.072 2.795 2.716 0.615 0.657
Ours ($\tau$) 88.1% 0.958 0.704 0.626 3.099 2.928 2.852 0.532 0.615
[…]
Ours ($\tau$, rev) 40.8 % 0.749 0.548 0.487 2.970 2.743 2.659 0.640 0.685
[…]
@@
4.2 Results on Synthesizable Analog Generation
@@
In Table 1, we see our method outperforming SynNet across both dimensions of similarity (how “analog” compounds are) and diversity (how different the compounds are). Additionally, our method achieves lower SA Score, which is a proxy for synthetic accessibility that rewards simpler molecules. Guided by a set of simple yet expressive syntactic templates, our model simultaneously produces more diverse and structurally simple molecules without sacrificing one for the other. Additionally, our policy network is well-suited to navigate these simple yet horizon structures, enabling a $6\%$ higher reconstruction accuracy after training on the same dataset. Combining these three dimensions, we can conclude our method is the superior one for the task of synthesizable analog generation. To better understand which design choices are responsible for the performance, we provided a comprehensive analysis of the policy network in App. F. We begin in App. F.3 by elaborating on the main distinction of our method vs existing works, highlight the novelty of our formulation, and motivate an auxiliary training task that takes inspiration from cutting-edge ideas in inductive program synthesis. We then perform several key ablations in App. F.4, using concrete examples to highlight success and failure cases. Lastly, we perform a step-by-step walkthrough of our decoding algorithm in App. F.8, visualizing the evolution of attention weights to showcase the full-horizon awareness of our surrogate and the dynamic incorporation of new information. These analyses shed insights into why our surrogate works, and points to future extensions to make it even better.
Table 2: Our model’s average performance across 13 TDC oracles compared to baselines compiled in Gao et al. (2022). We limit to 1000 oracle calls each run and normalize oracle outputs to $[0,1]$. We report to mean score, AUC, and SA scores (Ertl & Schuffenhauer, 2009) of the top 10 molecules. Methods are categorized by Gao et al. (2022) and, for brevity, we display only the top three algorithms within each category with respect their AUC. The best method per column is bolded and the best synthesis-based method is underlined. See App. H for the full results and experiment details.
Score $\uparrow$ AUC $\uparrow$ SA $\downarrow$
Category Method Value Rank Value Rank Value Rank
Screening Screening 0.426 20 0.377 20 3.097 8
MolPAL 0.472 16 0.444 15 3.018 4
String REINVENT 0.697 2 0.607 2 3.415 9
REINVENT-SELFIES 0.682 3 0.578 4 3.791 15
STONED 0.609 8 0.555 6 5.550 24
7 rows omitted $\cdots$
Graph Graph-GA 0.701 1 0.601 3 3.982 17
GPBO 0.642 6 0.570 5 3.954 16
DST 0.555 10 0.479 11 4.146 20
7 rows omitted $\cdots$
Synthesis SynNet 0.578 9 0.545 7 3.075 6
DoG-Gen 0.634 7 0.511 9 2.793 2
DoG-AE 0.460 18 0.450 14 2.857 3